### v1

# An example of the Jantzen filtration of a D-module

Abstract.
We compute the Jantzen filtration of a $\mathcal{D}$-module on the flag variety of $\operatorname{SL}_{2}(\mathbb{C})$. At each step in the computation, we illustrate the $\mathfrak{sl}_{2}(\mathbb{C})$-module structure on global sections to give an algebraic picture of this geometric computation. We conclude by showing that the Jantzen filtration on the $\mathcal{D}$-module agrees with the algebraic Jantzen filtration on its global sections, demonstrating a famous theorem of Beilinson–Bernstein.
1. Introduction
Jantzen filtrations arise in many situations in representation theory. The Jantzen filtration of a Verma module over a semisimple Lie algebra provides information on characters (the Jantzen sum formula) [Jan79], and gives representation-theoretic significance to coefficients of Kazhdan–Lusztig polynomials (the Jantzen conjectures) [BB93]. The Jantzen filtration of a Weyl module over a reductive algebraic group of positive characteristic is a helpful tool in the notoriously difficult problem of determining irreducible characters [Jan79]. Jantzen filtrations also play a critical role in the unitary algorithm of [AvLTV20], which determines the irreducible unitary representations of a real reductive group.
Though the utility of Jantzen filtrations in applications is primarily algebraic (providing information about characters or multiplicities of representations), establishing deep properties of the Jantzen filtration usually requires a geometric incarnation due to Beilinson–Bernstein. In [BB93], Beilinson–Bernstein introduce a $\mathcal{D}$-module version of the Jantzen filtration. Their geometric filtration aligns with the algebraic Jantzen filtration under Beilinson–Bernstein localization. Beilinson–Bernstein identify their geometric Jantzen filtration with the weight filtration on the $\mathcal{D}$-module, which provides them with powerful geometric tools to analyze its structure. In particular, they deduce the Jantzen conjecture: the semi-simplicity of the layers of the Jantzen filtration, and the compatibility of the Jantzen filtration with natural morphisms between Verma modules.
The constructions in [BB93] require technical and deep machinery in the theory of $\mathcal{D}$-modules and weight filtrations, and as such, may not be easily accessible to a reader unfamiliar with this geometric approach to representation theory. However, the persistent utility of Beilinson–Bernstein’s results indicates that the geometric Jantzen filtration is a critical tool.
In our experience, it is often enlightening, insightful, and non-trivial to describe a difficult construction in a simple example. This is the purpose of this paper: to illustrate the construction of Beilinson–Bernstein in the simplest non-trivial example. We mean “illustrate” in the most literal sense – we include eight figures in which we draw pictures of the geometric objects arising in Beilinson–Bernstein’s construction. Our hope is that by giving a concrete visual description of this fundamental construction, we are able to illustrate its natural beauty and provide readers with intuition for the general construction.
1.1. Structure of the paper
The focus of this paper is an extended $\operatorname{SL}_{2}(\mathbb{C})$ example, and the majority of the exposition is dedicated to the computation of the geometric Jantzen filtration in this example. We begin the paper in §2 with a brief outline of the construction of the geometric Jantzen filtration for a general Lie algebra. In §3, we shift our focus to $\operatorname{SL}_{2}(\mathbb{C})$ for the remainder of the paper. Our first computation in §$\S 3.1$ establishes an algebra homomorphism from the extended universal enveloping algebra to global differential operators on base affine space. This algebra homomorphism is what allows us to view the global sections of $\mathcal{D}$-modules as modules over the (extended) universal enveloping algebra. In §3.2, we give some background on $H$-monodromic $\mathcal{D}$-modules, and explain their relationship to modules over twisted sheaves of differential operators. In §3.3, we introduce the $\mathcal{D}$-modules whose global sections contain Verma modules and dual Verma modules — these are the $\mathcal{D}$-modules which we will endow with geometric Jantzen filtrations. We illustrate the $\mathfrak{sl}_{2}(\mathbb{C})$-module structure on their global sections in Figures 1 and 2. In §3.4, we introduce the maximal extension functor, which gives us the deformation necessary for the construction of the Jantzen filtration. We compute the maximal extension of the structure sheaf on an open subset of base affine space, and illustrate in Figures 3 and 4 how deformed Verma modules and deformed dual Verma modules arise in this computation. We illustrate in Figures 5 and 6 the global sections of the maximal extension, identifying them with the big projective in category $\mathcal{O}$. In §3.5, we introduce monodromy filtrations in abelian categories, and define the geometric Jantzen filtration on a $\mathcal{D}$-module using this set-up. We compute the monodromy filtration of the maximal extension, and illustrate its global sections in Figure 7. This specializes to the geometric Jantzen filtration on certain sub- and quotient sheaves. We finish in §3.6 by relating the geometric Jantzen filtration to the algebraic one. This step is critical, and is given only minimal justification in [BB93]. We introduce the algebraic Jantzen filtration on a Verma module in §3.6.1, then explain why the global sections of the geometric Jantzen filtration align with the algebraic Jantzen filtration in §3.6.2. Figure 8 illustrates this relationship in our example.
2. The Jantzen filtration of a $\mathcal{D}$-module
This paper is concerned with the example of $\operatorname{SL}_{2}(\mathbb{C})$. However, some amount of general theory will be helpful to set the scene. In this section, we briefly orient the reader with the necessary general theory. More details of this construction can be found in [BB93].
2.1. The monodromy filtration
Geometric Jantzen filtrations are intimately related to monodromy filtrations. Given an object $A$ in an abelian category $\mathcal{A}$ and a nilpotent endomorphism $s\in\operatorname{End}(A)$, there is a unique filtration of $A$ satisfying certain symmetry properties with respect to $s$. (See §3.5 for a precise definition.) This is the monodromy filtration of $A$. It induces filtrations on the sub and quotient objects $\ker(s:A\rightarrow A)$ and $\operatorname{coker}(s:A\rightarrow A)$ in the natural way.
2.2. Geometric Jantzen filtration
Certain $\mathcal{D}$-modules come naturally equipped with nilpotent endomorphisms, and thus acquire monodromy filtrations. If $Y$ is a smooth algebraic variety with a fixed regular function $f:Y\rightarrow\mathbb{A}^{1}$, the maximal extension $\Xi_{f}\mathcal{M}_{U}$ of a holonomic $\mathcal{D}_{U}$-module $\mathcal{M}_{U}$ on $U=f^{-1}(\mathbb{A}^{1}-\{0\})$ is an object in the abelian category of holonomic $\mathcal{D}_{Y}$-modules which has a natural nilpotent endomorphism $s$. Hence it has a monodromy filtration. Moreover, by the construction of the maximal extension functor (which is outlined in §3.4), $\ker(s:\Xi_{f}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U})=j_{!}\mathcal{M}_{U}$ and $\operatorname{coker}(s:\Xi_{f}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U})=j_{+}\mathcal{M}_{U}$, where $j:U\hookrightarrow Y$ is inclusion. In this way, we obtain filtrations of the $\mathcal{D}_{Y}$-modules $j_{!}\mathcal{M}_{U}$ and $j_{+}\mathcal{M}_{U}$ from the monodromy filtration of $\Xi_{f}\mathcal{M}_{U}$. These are the geometric Jantzen filtrations.
2.3. Geometric Jantzen filtration on Harish-Chandra sheaves
If $(\mathfrak{g},K)$ is a Harish-Chandra pair [^1] Meaning that $\mathfrak{g}$ is a complex semisimple Lie algebra and $K$ is an algebraic group equipped with a morphism $K\rightarrow\mathrm{Int}\mathfrak{g}$ whose differential is injective, with the property that $K$ acts with finitely many orbits on the flag variety of $\mathfrak{g}$. with admissible $K$-orbits [^2] See [BB93, 3.4.1(i)] for this rather technical definition. In the cases where $K=N$ is a maximal nilpotent subgroup or $(\mathfrak{g},K)$ is a symmetric pair, all orbits are admissible. in the flag variety $X$ of $\mathfrak{g}$, then the !!!- and $+$-standard $\mathcal{D}_{X}$-modules obtained by pushing forward $K$-equivariant connections on $K$-orbits can be made to fit into the framework of §2.2, and thus acquire geometric Jantzen filtrations. To explain more details of this construction we must introduce more notation.
Let $G$ be the simply connected semisimple Lie group associated to $\mathfrak{g}$, and fix a Borel subgroup $B\subset G$. Let $N\subset B$ denote the unipotent radical of $B$, and set $H:=B/N$ to be the abstract maximal torus of $G$ [CG97, Lem. 6.1.1], with Lie algebra $\mathfrak{h}:=\operatorname{Lie}H$. Let $\widetilde{X}:=G/N$ be the base affine space corresponding to $\mathfrak{g}$, and identify $X$ with $G/B$. Let $\pi:\widetilde{X}\rightarrow X$ be projection. This is a principal $G$-equivariant $H$-bundle with respect to the right action of $H$ on $\widetilde{X}$ by right multiplication.
For a $K$-orbit $Q$ in $X$, denote by $\widetilde{Q}=\pi^{-1}(Q)$ the corresponding union of $K$-orbits in $\widetilde{X}$. A choice of an integral weight $\lambda\in\mathfrak{h}^{*}$ (the “deformation direction”) determines a function $f_{\lambda}:\overline{\widetilde{Q}}\rightarrow\mathbb{A}^{1}$ on the closure of $\widetilde{Q}$ such that $f_{\lambda}^{-1}(\mathbb{A}^{1}-\{0\})=\widetilde{Q}$. This function extends to a regular function on $\widetilde{X}$, which determines a maximal extension functor $\Xi_{f_{\lambda}}:\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{U})\rightarrow\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\widetilde{X}})$, where $U$ is the preimage in $\widetilde{X}$ of $\mathbb{A}^{1}-\{0\}$ under the extension of $f_{\lambda}$, as in §2.2. Restricting $\Xi_{f_{\lambda}}$ to the category of holonomic $\mathcal{D}_{U}$-modules supported on $\widetilde{Q}$ results in a functor
$\Xi_{f_{\lambda}}:\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\widetilde{Q}})\rightarrow\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\overline{\widetilde{Q}}}).$
Let $\mathcal{M}_{\widetilde{Q}}$ be an irreducible $K$-equivariant connection on $\widetilde{Q}$. Using the procedure outlined above, the modules $j_{\widetilde{Q}!}\mathcal{M}_{\widetilde{Q}}$ and $j_{\widetilde{Q}+}\mathcal{M}_{\widetilde{Q}}$ acquire from $\Xi_{f_{\lambda}}\mathcal{M}_{\widetilde{Q}}$ geometric Jantzen filtrations. Here $j_{\widetilde{Q}}:\widetilde{Q}\hookrightarrow\overline{\widetilde{Q}}$ is inclusion of $\widetilde{Q}$ into its closure. Because $\overline{\widetilde{Q}}$ is closed in $\widetilde{X}$, we can use Kashiwara’s theorem to lift these filtrations to filtrations of the standard $K$-equivariant $\mathcal{D}_{\widetilde{X}}$-modules $i_{\widetilde{Q}!}\mathcal{M}_{\widetilde{Q}}$ and $i_{\widetilde{Q}+}\mathcal{M}_{\widetilde{Q}}$, for $i_{\widetilde{Q}}:\widetilde{Q}\hookrightarrow\widetilde{X}$ the inclusion.
2.4. Relationship to algebraic Jantzen filtration
Classical results in Lie theory [Bei81, Milb] relate global sections of standard $K$-equivariant $\mathcal{D}_{X}$-modules to well-known classes of $\mathfrak{g}$-modules, such as Verma modules or principal series representations, depending on specific choices of $K$. These $\mathfrak{g}$-modules also admit algebraic Jantzen filtrations (see §3.6.1), which are usually defined in terms of contravariant forms. Under Beilinson–Bernstein localisation, the algebraic and geometric Jantzen filtrations align. In §3.6.2 we explain this relationship in our example.
3. Example
Now we proceed with our example. For the remainder of this paper, set $G=\operatorname{SL}_{2}(\mathbb{C})$, and fix subgroups
$B=\Set{\begin{pmatrix}a&b\\ 0&a^{-1}\end{pmatrix}}{\,a\in\mathbb{C}^{*},\,b\in\mathbb{C}},\quad N=\Set{\begin{pmatrix}1&b\\ 0&1\end{pmatrix}}{b\in\mathbb{C}}$
and
$H=\Set{\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}}{a\in\mathbb{C}^{*}}.$
Let $\mathfrak{g}$, $\mathfrak{b}$, $\mathfrak{n}$, and $\mathfrak{h}$ be the corresponding Lie algebras, and $\bar{\mathfrak{n}}$ the opposite nilpotent subalgebra to $\mathfrak{n}$. Denote by
(3.1) $e=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\quad f=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\quad\text{and}\quad h=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
the standard basis elements of $\mathfrak{g}$, so $\mathfrak{n}=\mathbb{C}e$, $\mathfrak{h}=\mathbb{C}h$ and $\bar{\mathfrak{n}}=\mathbb{C}f$. Denote by $\mathcal{Z}(\mathfrak{g})$ the center of the universal enveloping algebra $\mathcal{U}(\mathfrak{g})$. The algebra $\mathcal{Z}(\mathfrak{g})$ is generated by the Casimir element
(3.2) $\Omega=h^{2}+2ef+2fe.$
Let
(3.3) $\gamma_{\mathrm{HC}}:\mathcal{U}(\mathfrak{g})\rightarrow\mathcal{U}(\mathfrak{h})$
be the projection onto the first coordinate of the direct sum decomposition
$\mathcal{U}(\mathfrak{g})=\mathcal{U}(\mathfrak{h})\oplus(\bar{\mathfrak{n}}\mathcal{U}(\mathfrak{g})+\mathcal{U}(\mathfrak{g})\mathfrak{n}).$
The restriction of $\gamma_{\mathrm{HC}}$ to $\mathcal{Z}(\mathfrak{g})$ is an algebra homomorphism.
Set $X=G/B$ and $\widetilde{X}=G/N$. Then $X$ is the flag variety of $\mathfrak{g}$, and we refer to $\widetilde{X}$ as base affine space. We identify $X$ with the complex projective line $\mathbb{C}\mathbb{P}^{1}$ via
(3.4) $\begin{pmatrix}x_{1}&*\\ x_{2}&*\end{pmatrix}B\mapsto(x_{1}:x_{2}),$
and $\widetilde{X}$ with $\mathbb{C}^{2}\backslash\{(0,0)\}$ via
(3.5) $\begin{pmatrix}x_{1}&*\\ x_{2}&*\end{pmatrix}N\mapsto(x_{1},x_{2}).$
There are left actions of $G$ on $X$ and $\widetilde{X}$ by left multiplication. Under the identifications (3.4) and (3.5), these actions are given by
(3.6) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot(x_{1}:x_{2})=(ax_{1}+bx_{2}:cx_{1}+dx_{2})$
and
(3.7) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot(x_{1},x_{2})=(ax_{1}+bx_{2},cx_{1}+dx_{2}).$
Because $H$ normalizes $N$, there is also a right action of $H$ on $G/N$ by right multiplication. Under the identification (3.5), this action is given by
(3.8) $(x_{1},x_{2})\cdot\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}=(ax_{1},ax_{2}).$
The natural $G$-equivariant quotient map
(3.9) $\pi:\widetilde{X}\rightarrow X$
is an $H$-torsor over $X$. In the language of [BB93, §2.5], this provides an “$H$-monodromic structure” on $X$.
For an algebraic variety $Y$, we denote by $\mathcal{O}_{Y}$ the structure sheaf on $Y$, and by $\mathcal{O}(Y)=\Gamma(Y,\mathcal{O}_{Y})$ the algebra of global regular functions. We denote by $\mathcal{D}_{Y}$ the sheaf of differential operators on $Y$, and $\mathcal{D}(Y)=\Gamma(Y,\mathcal{D}_{Y})$ the global differential operators.
Base affine space $\widetilde{X}$ is a quasi-affine variety, with affine closure $\overline{\widetilde{X}}=\mathbb{A}^{2}$. Throughout this text, we will make use the following facts about quasi-affine varieties. Let $Y$ be an irreducible quasi-affine variety, openly embedded in an affine variety $\overline{Y}$.
  • If $Y$ is normal with $\mathrm{codim}_{\overline{Y}}(\overline{Y}\backslash Y)\geq 2$, then $\mathcal{O}(Y)=\mathcal{O}(\overline{Y})$ and $\mathcal{D}(Y)=\mathcal{D}(\overline{Y})$. (In particular, for $Y=\widetilde{X}$, this implies that global differential operators are nothing more than the Weyl algebra in 2 variables [^3] Outside of $\mathfrak{sl}_{2}(\mathbb{C})$, the situation is less straightforward. For a Lie algebra $\mathfrak{g}\neq\mathfrak{sl}_{2}(\mathbb{C})^{m}$, the affine closure of the corresponding base affine space is singular, and the ring of global differential operators can be quite complicated, see, for example, [LS06]..)
  • The variety $\overline{Y}$ is $D$-affine, meaning that the global sections functor induces an equivalence of categories between the category of quasi-coherent $\mathcal{D}_{\overline{Y}}$-modules and the category of modules over $\mathcal{D}(\overline{Y})$.
  • Since the inclusion $i:Y\rightarrow\overline{Y}$ is an open immersion, the restriction functor $i^{+}$ on the corresponding categories of $\mathcal{D}$-modules is exact, and commutes with pushforwards from open affine subvarieties.
The facts listed above allow us to move freely between $\mathcal{D}_{\widetilde{X}}$-modules and $\mathcal{D}(\mathbb{A}^{2})$-modules. We will do this periodically in computations.
3.1. The map $\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{Z}(\mathfrak{g})}\mathcal{U}(\mathfrak{h})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})$
Our strategy for gaining intuition about the $\mathcal{D}_{\widetilde{X}}$-modules arising in the construction of the Jantzen filtration is to illustrate the $\mathfrak{g}$-module structure on their global sections. This will give us an algebraic snapshot as to what is happening at each step in the construction sketched in Section 2. The first step is to differentiate the actions (3.7) and (3.8) to obtain a map $\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{Z}(\mathfrak{g})}\mathcal{U}(\mathfrak{h})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})$. This map provides the $\mathfrak{g}$-module structure on the global sections of $\mathcal{D}_{\widetilde{X}}$-modules. We dedicate this section to the computation of this map.
By differentiating the left action of $G$ in (3.7), we obtain an algebra homomorphism
(3.10) $L:\mathcal{U}(\mathfrak{g})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}}),\hskip 5.69054ptg\mapsto L_{g}$
given by the formula
(3.11) $L_{g}f(x)=\frac{d}{dt}\bigg{|}_{t=0}f(\mathrm{exp}(tg)^{-1}x)$
for $g\in G$, $f\in\Gamma(\widetilde{X},\mathcal{O}_{\widetilde{X}})$, $x\in\widetilde{X}$. Computing the image of the basis (3.1) under the homomorphism (3.10) is straighforward. For example, the image of $e$ is given by the following computation using (3.7).
$e\cdot f(x_{1},x_{2})$ $=\frac{d}{dt}\bigg{|}_{t=0}f\left(\begin{pmatrix}1&-t\\ 0&1\end{pmatrix}\cdot(x_{1},x_{2})\right)$
$=\frac{d}{dt}\bigg{|}_{t=0}f(x_{1}-tx_{2},x_{2})$
$=-x_{2}\partial_{1}f(x_{1},x_{2}).$
Similar computations determine the image of $f$ and $h$:
(3.12) $L_{e}=-x_{2}\partial_{1},\quad L_{f}=-x_{1}\partial_{2},\quad L_{h}=-x_{1}\partial_{1}+x_{2}\partial_{2}.$
It is also useful to compute the image of the Casimir element (3.2) under the homomorphism $L$:
(3.13) $L_{\Omega}=x_{1}^{2}\partial_{1}^{2}+3x_{1}\partial_{1}+3x_{2}\partial_{2}+x_{2}^{2}\partial_{2}^{2}+2x_{1}x_{2}\partial_{1}\partial_{2}.$
Similarly, the right action of $H$ determines an algebra homomorphism
(3.14) $R:\mathcal{U}(\mathfrak{h})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}}),\quad g\mapsto R_{g}.$
Under this homomorphism, $h$ is sent to the Euler operator
(3.15) $R_{h}=x_{1}\partial_{1}+x_{2}\partial_{2}.$
Combining the homomorphisms $L$ (3.10) and $R$ (3.14), we obtain an algebra homomorphism
(3.16) $\mathcal{U}(\mathfrak{g})\otimes_{\mathbb{C}}\mathcal{U}(\mathfrak{h})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}});\quad g\otimes g^{\prime}\mapsto L_{g}R_{g^{\prime}}$
Lemma 3.1.
The homomorphism (3.16) factors through the quotient
(3.17) $\widetilde{\mathcal{U}}:=\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{Z}(\mathfrak{g})}\mathcal{U}(\mathfrak{h}),$
where $\mathcal{Z}(\mathfrak{g})$ acts on $\mathcal{U}(\mathfrak{h})$ via the Harish-Chandra projection $\gamma_{\mathrm{HC}}$ (3.3).
Proof.
Direct computation shows that the image of $\Omega\otimes 1$ and $1\otimes\gamma_{\mathrm{HC}}(\Omega)$ agree. Indeed,
$1\otimes\gamma_{\mathrm{HC}}(\Omega)=1\otimes(h^{2}+2h)\mapsto$ $R_{h}^{2}+2R_{h}$
$=(x_{1}\partial_{1}+x_{2}\partial_{2})^{2}+2(x_{1}\partial_{1}+x_{2}\partial_{2})$
$=x_{1}^{2}\partial_{1}^{2}+3x_{1}\partial_{1}+2x_{1}x_{2}\partial_{1}\partial_{2}+x_{2}^{2}\partial_{x}^{2}+3x_{2}\partial_{x}$
$=L_{\Omega}.$
We refer to the algebra $\widetilde{\mathcal{U}}$ as the extended universal enveloping algebra. By Lemma 3.1, we have an algebra homomorphism
(3.18) $\alpha:\widetilde{\mathcal{U}}\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}});\quad g\otimes g^{\prime}\mapsto L_{g}R_{g^{\prime}}.$
Global sections of $\mathcal{D}_{\widetilde{X}}$-modules have the structure of $\widetilde{\mathcal{U}}$-modules via $\alpha.$
3.2. Monodromic $\mathcal{D}_{X}$-modules
The $\mathcal{D}$-modules which play a role in our story have an additional structure: they are “$H$-monodromic”. It is necessary for our purposes to work with $H$-monodromic $\mathcal{D}$-modules on base affine space instead of $\mathcal{D}$-modules on the flag variety. This is due to the fact that the $\mathfrak{g}$-modules in the construction of the Jantzen filtration have generalized infinitesimal character, so they do not arise as global sections of modules over twisted sheaves of differential operators on the flag variety.
The machinery of $H$-monodromic $\mathcal{D}$-modules is rather technical, and the details of the construction are not strictly necessary for our computation of the Jantzen filtration below. However, we thought that it would be useful to describe this construction in a specific example to illustrate that the equivalences established in [BB93, §2.5] are quite clear for $\mathfrak{sl}_{2}(\mathbb{C})$. In this section, we describe the construction of $H$-monodromic $\mathcal{D}$-modules for $\mathfrak{sl}_{2}(\mathbb{\mathbb{C}})$ and explain how it relates to representations of Lie algebras. More details on the general construction can be found in [BB93, BG99].
Definition 3.2.
An $H$-monodromic $\mathcal{D}_{X}$-module is a weakly $H$-equivariant [^4] For a reference on weakly equivariant $\mathcal{D}$-modules, see [MP98, §4]. $\mathcal{D}_{\widetilde{X}}$-module.
There is an equivalent characterization of monodromic $\mathcal{D}_{X}$-modules in terms of $H$-invariant differential operators which is established in [BB93, §2.5.2]. This perspective makes the structures of our examples more transparent, so we will take this approach to monodromicity. Below we describe the construction for $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{C})$.
The right $H$-action in (3.8) induces a left $H$-action on $\mathcal{O}_{\widetilde{X}}$ and $\mathcal{D}_{\widetilde{X}}$. The $H$-action on $\mathcal{D}_{\widetilde{X}}$ satisfies the following relation: for $g\in H$, $\theta\in\mathcal{D}_{\widetilde{X}}$, and $f\in\mathcal{O}_{\widetilde{X}}$,
(3.19) $(g\cdot\theta)(g\cdot f)=g\cdot(\theta(f)).$
The $H$-action on $\mathcal{D}_{\widetilde{X}}$ induces an $H$-action on the sheaf $\pi_{*}(\mathcal{D}_{\widetilde{X}})$ by algebra automorphisms, where $\pi:\widetilde{X}\rightarrow X$ is the quotient map (3.9). Here $\pi_{*}$ is the $\mathcal{O}$-module direct image. Denote the sheaf of $H$-invariant sections of $\pi_{*}\mathcal{D}_{\widetilde{X}}$ by
(3.20) $\widetilde{\mathcal{D}}:=[\pi_{*}\mathcal{D}_{\widetilde{X}}]^{H}.$
This is a sheaf of algebras on $X$. Explicitly, on an open set $U\subseteq X$,
(3.21) $\widetilde{\mathcal{D}}(U)=\mathcal{D}_{\widetilde{X}}(\pi^{-1}(U))^{H}.$
Note that because $\pi$ is an $H$-torsor, $\pi^{-1}(U)$ is $H$-stable for any set $U$, so this construction is well-defined.
Let $\mathcal{M}(\mathcal{D}_{\widetilde{X}},H)_{\mathrm{weak}}$ be the category of weakly $H$-equivariant $\mathcal{D}_{\widetilde{X}}$-modules, and $\mathcal{M}(\widetilde{\mathcal{D}})$ be the category of $\widetilde{\mathcal{D}}$-modules. By [BB93, §1.8.9, §2.5.2], there is an equivalence of categories
(3.22) $\mathcal{M}(\mathcal{D}_{\widetilde{X}},H)_{\mathrm{weak}}\simeq\mathcal{M}(\widetilde{\mathcal{D}}).$
Hence we can study monodromic $\mathcal{D}_{X}$-modules by instead considering $\widetilde{\mathcal{D}}$-modules. For the remainder of the paper, we will take this to be our definition of monodromicity.
Definition 3.3.
An $H$-monodromic $\mathcal{D}_{X}$-module is a $\widetilde{\mathcal{D}}$-module, where $\widetilde{\mathcal{D}}$ is as in (3.20).
Remark 3.4.
(Relationship to twisted differential operators) The sheaf $\widetilde{\mathcal{D}}$ is a sheaf of $S(\mathfrak{h})$-algebras. In our example, the $S(\mathfrak{h})$-action is given by multiplication by the operator $R_{h}$ (3.15). In particular, we can consider $S(\mathfrak{h})$ as a subsheaf of $\widetilde{\mathcal{D}}$. In fact, it is the center [BB93, §2.5]. For $\lambda\in\mathfrak{h}^{*}$, denote by $\mathfrak{m}_{\lambda}\subset S(\mathfrak{h})$ the corresponding maximal ideal. The sheaf $\mathcal{D}_{\lambda}:=\widetilde{\mathcal{D}}/\mathfrak{m}_{\lambda}\widetilde{\mathcal{D}}$ is a twisted sheaf of differential operators (TDO) on $X$. Hence $\widetilde{\mathcal{D}}$-modules on which $R_{h}$ acts by eigenvalue $\lambda$ can be naturally identified with modules over the TDO $\mathcal{D}_{\lambda}$.
Modules over $\widetilde{\mathcal{D}}$ are directly related to modules over the extended universal enveloping algebra (3.17) via the global sections functor. The relationship is as follows. Because the left $G$-action and right $H$-action commute, the differential operators $L_{e},L_{f},$ and $L_{h}$ in (3.12) are $H$-invariant [^5] This can also be shown via direct computation using (3.12) and (3.15).. Hence the image of the homomorphism (3.18) is contained in $H$-invariant differential operators:
$\alpha(\widetilde{\mathcal{U}})\subseteq\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})^{H}.$
Composing $\alpha$ with $\Gamma(\pi_{*})$, we obtain a map
(3.23) $\widetilde{\mathcal{U}}\rightarrow\Gamma(X,\widetilde{\mathcal{D}}).$
Theorem 3.5.
[BB93, Lemma 3.2.2] The map (3.23) is an isomorphism.
Proof.
The theorem holds for general $\mathfrak{g}$. We will prove the theorem for $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{C})$ by direct computation.
We start by describing the sheaves $\pi_{*}\mathcal{D}_{\widetilde{X}}$ and $\widetilde{\mathcal{D}}$ on $X=\mathbb{C}\mathbb{P}^{1}$ by describing them on the open patches
(3.24) $U_{1}:=\mathbb{C}\mathbb{P}^{1}\backslash\{(0:1)\}\quad\text{and}\quad U_{2}:=\mathbb{C}\mathbb{P}^{1}\backslash\{(1:0)\}$
and giving gluing conditions. Set
(3.25) $V_{1}:=\pi^{-1}(U_{1})=\mathbb{C}^{2}\backslash V(x_{1})\quad\text{and}\quad V_{2}:=\pi^{-1}(U_{2})=\mathbb{C}^{2}\backslash V(x_{2}),$
where $V(f(x_{1},x_{2}))$ denotes the vanishing of the polynomial $f(x_{1},x_{2})$. By definition, we have
(3.26) $\pi_{*}\mathcal{D}_{\widetilde{X}}(U_{1})$ $=\mathcal{D}_{\widetilde{X}}(V_{1})=\mathcal{D}(\mathbb{C}^{2})[x_{1}^{-1}],$
(3.27) $\pi_{*}\mathcal{D}_{\widetilde{X}}(U_{2})$ $=\mathcal{D}_{\widetilde{X}}(V_{2})=\mathcal{D}(\mathbb{C}^{2})[x_{2}^{-1}].$
with obvious gluing conditions.
Using (3.19), we conclude that the $H$-action on $\mathcal{D}_{\widetilde{X}}$ is given by the local formulas
(3.28) $g\cdot x_{i}=gx_{i}\quad\text{and}\quad g\cdot\partial_{i}=g^{-1}\partial_{i},$
where $g\in H$ is regarded as an element of $\mathbb{C}^{\times}$ under the identification
$H\simeq\mathbb{C}^{\times},\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}\mapsto a.$
From this we obtain a local description of $\widetilde{\mathcal{D}}$, using (3.21):
(3.29) $\widetilde{\mathcal{D}}(U_{1})$ $=\langle x_{1}^{-1}x_{2},x_{1}\partial_{1},x_{1}\partial_{2},x_{2}\partial_{1},x_{2}\partial_{2}\rangle\subseteq\mathcal{D}_{\widetilde{X}}(V_{1})$
(3.30) $\widetilde{\mathcal{D}}(U_{2})$ $=\langle x_{1}x_{2}^{-1},x_{1}\partial_{1},x_{1}\partial_{2},x_{2}\partial_{1},x_{2}\partial_{2}\rangle\subseteq\mathcal{D}_{\widetilde{X}}(V_{2}).$
Hence the global sections are given by
(3.31) $\Gamma(X,\widetilde{\mathcal{D}})\simeq\langle x_{1}\partial_{1},x_{1}\partial_{2},x_{2}\partial_{1},x_{2}\partial_{2}\rangle=\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})^{H}\subseteq\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}}).$
Now, it is clear that
(3.32) $L_{e}=-x_{2}\partial_{1},\quad L_{f}=-x_{1}\partial_{2},\quad L_{h}+R_{h}=2x_{2}\partial_{2}\quad\text{and}\quad L_{h}-R_{h}=-2x_{1}\partial_{1}$
so the operators $L_{e},L_{f},L_{h}$ and $R_{h}$ generate $\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})^{H}$. Since $e\otimes 1,f\otimes 1,h\otimes 1$ and $1\otimes h$ generate $\widetilde{\mathcal{U}}$, this shows that the map (3.23) is surjective. Direct computations establish that
$[L_{e},L_{f}]=x_{2}\partial_{1}x_{1}\partial_{2}-x_{1}\partial_{2}x_{2}\partial_{1}=x_{2}\partial_{2}-x_{1}\partial_{1}=L_{h},$
$[L_{e},L_{h}]=x_{2}\partial_{1}(x_{1}\partial_{1}-x_{2}\partial_{2})-(x_{1}\partial_{1}-x_{2}\partial_{2})x_{2}\partial_{1}=2x_{2}\partial_{1}=-2L_{e},$
$[L_{f},L_{h}]=x_{1}\partial_{2}(x_{1}\partial_{1}-x_{2}\partial_{2})-(x_{1}\partial_{1}-x_{2}\partial_{2})x_{1}\partial_{2}=-2x_{1}\partial_{2}=2L_{f},$
$[L_{e},R_{h}]=[L_{f},R_{h}]=[L_{h},R_{h}]=0.$
Combining these computations with the fact that $L_{e},L_{f},L_{h}$ and $R_{h}$ are linearly independent shows that the relations satisfied by $L_{e},L_{f},L_{h},$ and $R_{h}$ are precisely those satisfied by $e\otimes 1,f\otimes 1,h\otimes 1$ and $1\otimes h.$ Therefore, the map (3.23) is also injective. ∎
The relationships described in this section can be summarised with the following commuting diagrams.
[…]
The composition of the top two arrows and the right-most arrow is the equivalence (3.22) (See [BB93, §1.8.9, §2.5.3] for more details.)
3.3. Verma modules and dual Verma modules
Using the map (3.18) constructed in Section 3.1, we can describe the $\widetilde{\mathcal{U}}$-module structure on various classes of $\mathcal{D}_{\widetilde{X}}$-modules. We will start by examining the $\mathcal{D}_{\widetilde{X}}$-modules $j_{+}\mathcal{O}_{U}$ and $j_{!}\mathcal{O}_{U}$, where $j:U\hookrightarrow\widetilde{X}$ is inclusion of the open $B$-orbit
(3.34) $U:=\mathbb{C}^{2}\backslash V(x_{2}).$
Here the $+$ and !!! indicate the $\mathcal{D}$-module push-forward functors, see [Mila]. These are the $\mathcal{D}_{\widetilde{X}}$-modules which will eventually be endowed with geometric Jantzen filtrations in Section 3.5. In this section, we describe the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})$ and $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$.
Because $j$ is an open embedding, the $\mathcal{D}_{\widetilde{X}}$-module $j_{+}\mathcal{O}_{U}$ is just the sheaf $\mathcal{O}_{U}$ with $\mathcal{D}_{\widetilde{X}}$-module structure given by the restriction of $\mathcal{D}_{U}$ to $\mathcal{D}_{\widetilde{X}}\subseteq\mathcal{D}_{U}$. Hence the global sections of $j_{+}\mathcal{O}_{U}$ can be identified with the ring
(3.35) $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})=\mathbb{C}[x_{1},x_{2},x_{2}^{-1}].$
The operators $L_{e},L_{f},L_{h}$, and $R_{h}$ from (3.12) and (3.15) act on monomials $x_{1}^{m}x_{2}^{n}$ for $m\geq 0$ $n\in\mathbb{Z}$ by the formulas
(3.36) $L_{e}\cdot x_{1}^{m}x_{2}^{n}$ $=-mx_{1}^{m-1}x_{2}^{n+1},$
(3.37) $L_{f}\cdot x_{1}^{m}x_{2}^{n}$ $=-nx_{1}^{m+1}x_{2}^{n-1},$
(3.38) $L_{h}\cdot x_{1}^{m}x_{2}^{n}$ $=(n-m)x_{1}^{m}x_{2}^{n},$
(3.39) $R_{h}\cdot x_{1}^{m}x_{2}^{n}$ $=(m+n)x_{1}^{m}x_{2}^{n}.$
Using (3.36)-(3.39), we can illustrate the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})$ using nodes and colored arrows. We do this in Figure 1. The monomials $x_{1}^{m}x_{2}^{n}$ for $m\in\mathbb{Z}_{\geq 0}$ and $n\in\mathbb{Z}$ form a basis for $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})$. The green arrows illustrate the action of the operator $L_{e}$ on basis elements, the red arrows the action of $L_{f}$, and the blue arrows the action of $L_{h}$. If an operator acts by zero, no arrow is included. The $R_{h}$-eigenspaces are highlighted in grey, with corresponding eigenvalues listed below.
Figure 1. Dual Verma modules arise as global sections of $j_{+}\mathcal{O}_{U}$.
Remark 3.6.
We make the following observations about the $\mathcal{D}_{\widetilde{X}}$-module $j_{+}\mathcal{O}_{U}$ and its global sections.
  1. (1)
    As a $\widetilde{\mathcal{U}}$-module, $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})$ decomposes into a direct sum of submodules, each of which is an $R_{h}$-eigenspace corresponding to an integer eigenvalue:
    $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})=\bigoplus_{n\in\mathbb{Z}}\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})_{n}$
    In Figure 1, these eigenspaces are highlighted in grey.
  2. (2)
    As a $\mathcal{U}(\mathfrak{g})$-module, the $R_{h}$-eigenspace $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})_{n}$ of eigenvalue $n$ is isomorphic to the dual Verma module of highest weight $n$. In particular, it is irreducible if $n<0$, and it has a unique irreducible finite-dimensional submodule if $n\geq 0$.
  3. (3)
    The sheaf $\pi_{*}j_{+}\mathcal{O}_{U}$ is a monodromic $\mathcal{D}_{X}$-module because it admits an action of $\widetilde{\mathcal{D}}$ (Definition 3.3). For each positive integer $n$, $\pi_{*}j_{+}\mathcal{O}_{U}$ has a subsheaf $(\pi_{*}j_{+}\mathcal{O}_{U})_{n}$ on which $R_{h}$ acts locally by the eigenvalue $n$. These subsheaves are $\mathcal{D}_{n}$-modules, where $\mathcal{D}_{n}$ is the twisted sheaf of differential operators as defined in Remark 3.4. These are exactly the $\mathcal{D}_{n}$-modules appearing in [Rom21, §6, Fig. 4].
Next we will describe $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{\widetilde{X}})$. This is slightly more involved. By definition,
(3.40) $j_{!}=\mathbb{D}_{\widetilde{X}}\circ j_{+}\circ\mathbb{D}_{U},$
where $\mathbb{D}$ denotes the holonomic duality functor. Explicitly, for a smooth algebraic variety $Y$ and a holonomic $\mathcal{D}_{Y}$-module $\mathcal{V}$,
(3.41) $\mathbb{D}_{Y}(\mathcal{V}):=\operatorname{Ext}_{\mathcal{D}_{Y}}^{\dim Y}(\mathcal{V},\mathcal{D}_{X}).$
This is a well-defined functor from the category of holonomic $\mathcal{D}_{Y}$-modules to itself [HTT08, Cor. 2.6.8].
The first two steps of the composition (3.40) are straightforward to compute. The right $\mathcal{D}_{U}$-module $\mathbb{D}_{U}\mathcal{O}_{U}$ is just the sheaf $\mathcal{O}_{U}$, viewed as a right $\mathcal{D}_{U}$-module via the natural right action. Then, since $j$ is an open immersion, $j_{+}\mathbb{D}_{U}\mathcal{O}_{U}$ is the sheaf $\mathcal{O}_{U}$ with right $\mathcal{D}_{\widetilde{X}}$-module structure given by restriction to $\mathcal{D}_{\widetilde{X}}\subset\mathcal{D}_{U}$.
To apply $\mathbb{D}_{\widetilde{X}}$ to $j_{+}\mathbb{D}_{U}\mathcal{O}_{U}$, we must take a projective resolution of $j_{+}\mathbb{D}_{U}\mathcal{O}_{U}$. First, we make the identification
(3.42) $j_{+}\mathbb{D}_{U}\mathcal{O}_{U}\simeq\langle\partial_{1},\partial_{2}\rangle\mathcal{D}_{U}\backslash\mathcal{D}_{U}.$
We take the following free (hence projective) resolution of $\langle\partial_{1},\partial_{2}\rangle\mathcal{D}_{U}\backslash\mathcal{D}_{U}$:
(3.43) $00\leftarrow\langle\partial_{1},\partial_{2}\rangle\mathcal{D}_{U}\backslash\mathcal{D}_{U}\xleftarrow{\epsilon}\mathcal{D}_{\widetilde{X}}\xleftarrow{d_{0}}\mathcal{D}_{\widetilde{X}}\oplus\mathcal{D}_{\widetilde{X}}\xleftarrow{d_{1}}\mathcal{D}_{\widetilde{X}}\xleftarrow{d_{2}}$ 0
where the maps are defined by
$\epsilon:1$ $\mapsto x_{2}^{-1},$
$d_{0}:(\theta_{1},\theta_{2})$ $\mapsto\partial_{1}\theta_{1}-x_{2}\partial_{2}\theta_{2},$
$d_{1}:1$ $\mapsto(x_{2}\partial_{2},\partial_{1}).$
Applying $\mathrm{Hom}_{\mathcal{D}_{\widetilde{X}},r}(-,\mathcal{D}_{\widetilde{X}})$ to this complex, we obtain the complex
[…]
where $d_{i}^{*}$ sends a morphism $f$ to $f\circ d_{i}$.
Because the module $j_{+}\mathbb{D}_{U}\mathcal{O}_{U}$ is holonomic, the complex (3.44) only has nonzero cohomology in degree $2$. This can also be seen by direct computation. By identifying $\mathrm{Hom}_{\mathcal{D}_{\widetilde{X}},r}(\mathcal{D}_{\widetilde{X}},\mathcal{D}_{\widetilde{X}})\simeq\mathcal{D}_{\widetilde{X}}$ via $f\mapsto f(1)$ and $\mathrm{Hom}_{\mathcal{D}_{\widetilde{X}},r}(\mathcal{D}_{\widetilde{X}}\oplus\mathcal{D}_{\widetilde{X}},\mathcal{D}_{\widetilde{X}})\simeq\mathcal{D}_{\widetilde{X}}\oplus\mathcal{D}_{\widetilde{X}}$ via $f\mapsto(f(1,0),f(0,1))$, we see that
$\ker d_{2}^{*}\simeq\mathcal{D}_{\widetilde{X}}\text{ and }\operatorname{im}d_{1}^{*}\simeq\mathcal{D}_{\widetilde{X}}\langle\partial_{1},x_{2}\partial_{2}\rangle.$
Hence,
(3.45) $j_{!}\mathcal{O}_{U}\simeq\mathcal{D}_{\widetilde{X}}/\mathcal{D}_{\widetilde{X}}\langle\partial_{1},x_{2}\partial_{2}\rangle.$
Now we can describe the global sections of $j_{!}\mathcal{O}_{U}$ and illustrate their $\widetilde{\mathcal{U}}$-module structure, as we did for $\mathcal{O}_{\widetilde{X}}$ and $j_{+}\mathcal{O}_{U}$. The monomials $x_{1}^{m}x_{2}^{n}$ and $x_{1}^{m}\partial_{2}^{n}$ for $m,n\geq 0$ form a basis for $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$. The action of $L_{e},L_{f},L_{h}$ and $R_{h}$ on $x^{m}_{1}x_{2}^{n}$ for $m\geq 0$ and $n>0$ is given by equations (3.36)-(3.39). The action of $L_{e},L_{f},L_{h}$ and $R_{h}$ on $x^{m}_{1}\partial_{2}^{n}$ for $m\geq 0$ and $n>0$ is given by
(3.46) $L_{e}\cdot x_{1}^{m}\partial_{2}^{n}$ $=m(n-1)x_{1}^{m-1}\partial_{2}^{n-1},$
(3.47) $L_{f}\cdot x_{1}^{m}\partial_{2}^{n}$ $=-x_{1}^{m+1}\partial_{2}^{n+1},$
(3.48) $L_{h}\cdot x_{1}^{m}\partial_{2}^{n}$ $=-(m+n)x_{1}^{m}\partial_{2}^{n},$
(3.49) $R_{h}\cdot x_{1}^{m}\partial_{2}^{n}$ $=(m-n)x_{1}^{m}\partial_{2}^{n}.$
The action of $L_{e}$ on $x_{1}^{m}$ is given by (3.36), the action of $L_{f}$ on $x_{1}^{m}$ is given by (3.47), and the actions of $L_{h}$ and $R_{h}$ on $x^{m}$ are given by either (3.38)-(3.39) or (3.48)-(3.49).
We illustrate the $\widetilde{\mathcal{U}}$-module structure of $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$ in Figure 2. The colors indicate the same operators as in the earlier example: green is $L_{e}$, red is $L_{f}$, blue is $L_{h}$, and $R_{h}$-eigenspaces are highlighted in grey, with corresponding eigenvalues listed below.
Figure 2. Verma modules arise as global sections of $j_{!}\mathcal{O}_{U}$.
Remark 3.7.
We make the following observations about $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$.
  1. (1)
    As a $\widetilde{\mathcal{U}}$-module, $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$ decomposes into a direct sum of submodules, each of which is an $R_{h}$-eigenspace corresponding to an integer eigenvalue. Again, these eigenspaces are highlighted in grey.
  2. (2)
    As a $\mathcal{U}(\mathfrak{g})$-module, the $R_{h}$-eigenspace of $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$ of eigenvalue $n$ is isomorphic to the Verma module of highest weight $n$. In particular, it is irreducible if $n<0$, and it has a unique irreducible finite-dimensional quotient if $n\geq 0$.
  3. (3)
    The sheaf $\pi_{*}j_{!}\mathcal{O}_{U}$ is a monodromic $\mathcal{D}_{X}$-module. For each positive integer $n$, $\pi_{*}j_{!}\mathcal{O}_{U}$ has a subsheaf $(\pi_{*}j_{!}\mathcal{O}_{U})_{n}$ on which $R_{h}$ acts locally by the eigenvalue $n$. These subsheaves are modules over the TDO $\mathcal{D}_{n}$ (Remark 3.4). These are exactly the $\mathcal{D}_{n}$-modules appearing in [Rom21, §6, Fig. 3].
Remark 3.8.
It is interesting to contrast the computations above to the computations in [Rom21]. By working over base affine space instead of the flag variety, we are able to study all integral Verma and dual Verma modules at once using $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$. This is different from the approach in [Rom21, §6], where Verma and dual Verma modules are studied one at a time by examining modules over the TDO $\mathcal{D}_{n}$ as $n$ varies. The relationship between these two approaches comes from Remark 3.4: one can view $\widetilde{\mathcal{D}}$ as a family of TDOs parametrized by $\mathfrak{h}^{*}$, and modules over $\widetilde{\mathcal{D}}$ can be viewed as families of modules over those TDOs.
3.4. The maximal extension $\Xi_{\rho}\mathcal{O}_{U}$
To describe the geometric Jantzen filtrations on the $\mathcal{D}_{\widetilde{X}}$-modules $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$, it is necessary to introduce the maximal extension functor
$\Xi_{\rho}:\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{U})\rightarrow\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\widetilde{X}}).$
This functor (defined in (3.55)) extends $j_{+}$ and $j_{!}$ (see (3.57) - (3.58)), so it is a natural way to study both modules $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$ at once. In this section, we give the construction of $\Xi_{\rho}$, then describe the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$.
To start, we recall the construction of maximal extension for $\mathcal{D}$-modules, which is a special case of the construction in [Bei87], which produces the maximal extension and nearby cycles functors. Let $Y$ be a smooth variety, $f:Y\rightarrow\mathbb{A}^{1}$ a regular function, and
[…]
the corresponding open-closed decomposition of $Y$. For $n\in\mathbb{N}$, denote by
(3.51) $I^{(n)}:=\left(\mathcal{O}_{\mathbb{A}^{1}-\{0\}}\otimes\mathbb{C}[s]/s^{n}\right)t^{s}$
the free rank $1$ $\mathcal{O}_{\mathbb{A}^{1}-\{0\}}\otimes\mathbb{C}[s]/s^{n}$-module generated by the symbol $t^{s}$. The action $\partial_{t}\cdot t^{s}=st^{-1}t^{s}$ gives $I^{(n)}$ the structure of a $\mathcal{D}_{\mathbb{A}^{1}-\{0\}}$-module. Any $\mathcal{D}_{U}$-module $\mathcal{M}_{U}$ can be deformed using $I^{(n)}$: set
(3.52) $f^{s}\mathcal{M}_{U}^{(n)}:=f^{+}I^{(n)}\otimes_{\mathcal{O}_{U}}\mathcal{M}_{U}$
to be the $\mathcal{D}_{U}$-module obtained by twisting $\mathcal{M}_{U}$ by $I^{(n)}$. Note that $f^{s}\mathcal{M}_{U}^{(1)}=\mathcal{M}_{U}$.
Assume that $\mathcal{M}_{U}$ is holonomic. Denote by
(3.53) $\mathrm{can}:j_{!}f^{s}\mathcal{M}_{U}^{(n)}\rightarrow j_{+}f^{s}\mathcal{M}_{U}^{(n)}$
the canonical map between !!! and $+$ pushforward, and
(3.54) $s^{1}(n):j_{!}f^{s}\mathcal{M}_{U}^{(n)}\rightarrow j_{+}f^{s}\mathcal{M}_{U}^{(n)}$
the composition of $\mathrm{can}$ with multiplication by $s$. For large enough $n$, the cokernel of $s^{1}(n)$ stabilizes; i.e., $\operatorname{coker}s^{1}(n)=\operatorname{coker}s^{1}(n+k)$ for all $k>0$. For $n\gg 0$, define the $\mathcal{D}_{Y}$-module
(3.55) $\Xi_{f}\mathcal{M}_{U}:=\operatorname{coker}s^{1}(n),$
called the maximal extension of $\mathcal{M}_{U}$. By construction, this module comes equipped with the nilpotent endomorphism $s$. The corresponding functor
(3.56) $\Xi_{f}:\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{U})\rightarrow\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{Y})$
is exact [BB93, Lemma 4.2.1(i)]. Moreover, there are canonical short exact sequences [BB93, Lemma 4.2.1 (ii)’]
(3.57) $0\rightarrow j_{!}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U}\rightarrow\operatorname{coker}(\mathrm{can})\rightarrow 0$
(3.58) $0\rightarrow\operatorname{coker}(\mathrm{can})\rightarrow\Xi_{f}\mathcal{M}_{U}\rightarrow j_{+}\mathcal{M}_{U}\rightarrow 0$
with $j_{!}=\ker(s:\Xi_{f}\rightarrow\Xi_{f})$ and $j_{+}=\operatorname{coker}(s:\Xi_{f}\rightarrow\Xi_{f})$.
Now, we apply this general construction in the setting of our example. Let $\widetilde{X}$ and $U$ be as above (see (3.5) and (3.34)), and let $f_{\rho}$ be the function [^6] This choice of function corresponds to the deformation direction $\rho\in\mathfrak{h}^{*}$, see Remark 3.10..
(3.59) $f_{\rho}:\widetilde{X}\rightarrow\mathbb{A}^{1};(x_{1},x_{2})\mapsto x_{2}.$
For a variety $Y$, set
(3.60) $\mathcal{A}_{Y}:=\mathcal{D}_{Y}\otimes\mathbb{C}[s]/s^{n}.$
We will compute the maximal extension $\Xi_{\rho}\mathcal{O}_{U}:=\Xi_{f_{\rho}}\mathcal{O}_{U}$ of the structure sheaf $\mathcal{O}_{U}$ using the construction above, then describe the $\widetilde{\mathcal{U}}$-module structure on its global sections. To clarify the exposition, we list each step as a subsection.
3.4.1. Step 1: Deformation
Let $I^{(n)}$ be as in (3.51). The deformed version of $\mathcal{O}_{U}$ is
(3.61) $f^{s}\mathcal{O}_{U}^{(n)}=f^{+}I^{(n)}=\mathcal{O}_{U}\otimes_{f^{-1}(\mathcal{O}_{\mathbb{A}^{1}-\{0\}})}f^{-1}(I^{(n)}).$
The global sections of $f^{s}\mathcal{O}_{U}^{(n)}$ are
(3.62) $(\mathbb{C}[x_{1},x_{2},x_{2}^{-1}]\otimes\mathbb{C}[s]/s^{n})t^{s},$
where the differentials $\partial_{1},\partial_{2}\in\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})$ act on the generator $t^{s}$ by
(3.63) $\partial_{1}\cdot t^{s}=0\text{ and }\partial_{2}\cdot t^{s}=sx_{2}^{-1}t^{s}.$
Alternatively, we can identify $f^{s}\mathcal{O}_{U}^{(n)}$ with a quotient of $\mathcal{A}_{U}$:
(3.64) $f^{s}\mathcal{O}_{U}^{(n)}=\mathcal{A}_{U}/\mathcal{A}_{U}\langle\partial_{1},x_{2}\partial_{2}-s\rangle.$
Both descriptions will be useful below.
3.4.2. Step 2: $+$-pushforward
Because $j:U\hookrightarrow\widetilde{X}$ is an open embedding, the $\mathcal{D}_{\widetilde{X}}$-module $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ is the sheaf $f^{s}\mathcal{O}_{U}^{(n)}$ with $\mathcal{D}_{\widetilde{X}}$-module structure given by restriction to $\mathcal{D}_{\widetilde{X}}\subset\mathcal{D}_{U}$. Under the identification (3.64), we have
(3.65) $j_{+}f^{s}\mathcal{O}_{U}^{(n)}=\mathcal{A}_{U}/\mathcal{A}_{U}\langle\partial_{1},x_{2}\partial_{2}-s\rangle,$
with $\mathcal{D}_{\widetilde{X}}$-action given by left multiplication.
It is interesting to examine the $\widetilde{\mathcal{U}}$-module structure on the global sections of this module. The operators $L_{e},L_{f},L_{h}$ and $R_{h}$ (3.12), (3.15) act on the monomial basis elements of (3.62) by the following formulas
(3.66) $L_{e}\cdot x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}$ $=-kx_{1}^{k-1}x_{2}^{\ell+1}s^{m}t^{s};$
(3.67) $L_{f}\cdot x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}$ $=(-s-\ell)x_{1}^{k+1}x_{2}^{\ell-1}s^{m}t^{s};$
(3.68) $L_{h}\cdot x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}$ $=(s-k+\ell)x_{1}^{k}x_{2}^{\ell}s^{m}t^{s};$
(3.69) $R_{h}\cdot x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}$ $=(s+k+\ell)x_{2}^{k}x_{2}^{\ell}s^{m}t^{s}.$
The resulting $\widetilde{\mathcal{U}}$-module has a natural filtration given by powers of $s$, and it decomposes into a direct sum of submodules spanned by monomials $\{x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}\}$ for fixed integers $k+\ell$. Each of these submodules has the structure of a deformed dual Verma module, as illustrated [^7] We omit the generator $t^{s}$ and the arrows corresponding to the $R_{h}$-action in Figure 3 for clarity. in Figure 3 for $k+\ell=0$.
Figure 3. Deformed dual Verma modules arise as global sections of $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$.
Moreover, one can compute that the Casimir element $L_{\Omega}$ (3.13) acts by
(3.70) $L_{\Omega}\cdot x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}=\left((k+\ell)^{2}+2(k+\ell)+2s(1+k+\ell)+s^{2}\right)x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}.$
Since $s$ is nilpotent, we can see from this computation that a high enough power of the operator
(3.71) $L_{\Omega}-\gamma_{\mathrm{HC}}(k+\ell)=2s(1+k+\ell)+s^{2}$
annihilates any monomial basis element. (Here $\gamma_{\mathrm{HC}}$ is the Harish-Chandra projection in (3.3).) Hence, the module $\Gamma(\widetilde{X},j_{+}f^{s}\mathcal{O}_{U}^{(n)})$ has generalised, but not strict, infinitesimal character.
3.4.3. Step 3: !!!-pushforward
Recall that $j_{!}=\mathbb{D}_{\widetilde{X}}\circ j_{+}\circ\mathbb{D}_{U}$, where $\mathbb{D}$ denotes holonomic duality, as in (3.41). We begin by computing the right $\mathcal{D}_{U}$-module $\mathbb{D}_{U}f^{s}\mathcal{O}_{U}^{(n)}$ by taking a projective resolution of $f^{s}\mathcal{O}_{U}^{(n)}$ as a left $\mathcal{A}_{U}$-module. This is straightforward using the description (3.62). The complex
(3.72) $0\xrightarrow{d_{2}}\mathcal{A}_{U}\xrightarrow{d_{1}}\mathcal{A}_{U}\oplus\mathcal{A}_{U}\xrightarrow{d_{0}}\mathcal{A}_{U}\xrightarrow{\epsilon}\mathcal{A}_{U}/\mathcal{A}_{U}\langle\partial_{1},x_{2}\partial_{2}-s\rangle\rightarrow 0$
where $\epsilon$ is the canonical quotient map, $d_{0}$ sends $(\theta_{1},\theta_{2})\in\mathcal{A}_{U}\oplus\mathcal{A}_{U}$ to $\theta_{1}\partial_{1}-\theta_{2}(x_{2}\partial_{2}-s)$, and $d_{1}$ sends $1\mapsto(x_{2}\partial_{2}-s,\partial_{1})$, is a free resolution of the left $\mathcal{A}_{U}$-module $f^{s}\mathcal{O}_{U}^{(n)}$. Applying the functor $\mathrm{Hom}_{\mathcal{A}_{U}}(-,\mathcal{A}_{U})$ and making the natural identification
(3.73) $\mathrm{Hom}_{\mathcal{A}_{U}}(\mathcal{A}_{U},\mathcal{A}_{U})\simeq\mathcal{A}_{U};\varphi\mapsto\varphi(1),$
of right $\mathcal{A}_{U}$-modules, we see that
(3.74) $\mathbb{D}_{U}f^{s}\mathcal{O}_{U}^{(n)}=\mathrm{Ext}_{\mathcal{A}_{U}}^{2}(f^{s}\mathcal{O}_{U}^{(n)},\mathcal{A}_{U})=\operatorname{im}d_{1}^{*}\backslash\ker d_{2}^{*}=\langle\partial_{1},x_{2}\partial_{2}-s\rangle\mathcal{A}_{U}\backslash\mathcal{A}_{U}.$
Here $d_{i}^{*}(\varphi)=\varphi\circ d_{i}$ for an appropriate homomorphism $\varphi$, and the right $\mathcal{A}_{U}$-module structure is given by right multiplication.
To finish the computation of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ we must take a projective resolution of this module. We do so following a similar process to the !!!-pushforward computation in Section 3.3. Denote by $I$ the right ideal $\langle\partial_{1},x_{2}\partial_{2}-s\rangle\mathcal{A}_{U}$ in $\mathcal{A}_{U}$. The complex
(3.75) $00\leftarrow I\backslash\mathcal{A}_{U}\xleftarrow{\epsilon}\mathcal{A}_{\widetilde{X}}\xleftarrow{d_{0}}\mathcal{A}_{\widetilde{X}}\oplus\mathcal{A}_{\widetilde{X}}\xleftarrow{d_{1}}\mathcal{A}_{\widetilde{X}}\xleftarrow{d_{2}}$ 0
with maps given by
(3.76) $\epsilon:1$ $\mapsto Ix_{2}^{-1};$
(3.77) $d_{0}:(\theta_{1},\theta_{2})$ $\mapsto x_{2}\partial_{1}\theta_{1}-(x_{2}^{2}\partial_{2}-x_{2}s)\theta_{2};$
(3.78) $d_{1}:1$ $\mapsto(x_{2}\partial_{2}-s,\partial_{1})$
is a free resolution of $\mathbb{D}_{U}f^{s}\mathcal{O}_{U}^{(n)}$ by right $\mathcal{A}_{\widetilde{X}}$-modules. Applying $\mathrm{Hom}_{\mathcal{A}_{\widetilde{X}},r}(-,\mathcal{A}_{\widetilde{X}})$ and making the natural identifications as above, we obtain
(3.79) $j_{!}f^{s}\mathcal{O}_{U}^{(n)}=\ker d_{2}^{*}/\operatorname{im}d_{1}^{*}=\mathcal{A}_{\widetilde{X}}/\mathcal{A}_{\widetilde{X}}\langle\partial_{1},x_{2}\partial_{2}-s\rangle.$
The left $\mathcal{A}_{\widetilde{X}}$-module structure is given by left multiplication.
Again, it is interesting to examine the $\widetilde{\mathcal{U}}$-module structure on the global sections of this module. The global sections of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ are spanned by monomials $x_{1}^{k}x_{2}^{\ell}s^{m}$ for $k,\ell\geq 0$ and $0\leq m<n$ and $x_{1}^{a}\partial_{2}^{b}s^{m}$ for $a,b\geq 0$ and $0\leq m<n$. For $\ell>0$, the $L_{e},L_{f},L_{h}$ and $R_{h}$-actions on the monomials $x_{1}^{k}x_{2}^{\ell}s^{m}$ are as in (3.66) - (3.69) (where we identify the generator $t^{s}$ of $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ with the coset containing $1$ in $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$), and the actions on the monomials $x_{1}^{a}\partial_{2}^{b}s^{m}$ are given by the following formulas:
(3.80) $L_{e}\cdot x_{1}^{a}\partial_{2}^{b}s^{m}$ $=a(b-1-s)x_{1}^{a-1}\partial_{2}^{b-1}s^{m};$
(3.81) $L_{f}\cdot x_{1}^{a}\partial_{2}^{b}s^{m}$ $=-x_{1}^{a+1}\partial_{2}^{b+1}s^{m};$
(3.82) $L_{h}\cdot x_{1}^{a}\partial_{2}^{b}s^{m}$ $=(s-a-b)x_{1}^{a}\partial_{2}^{b}s^{m};$
(3.83) $R_{h}\cdot x_{1}^{a}\partial_{2}^{b}s^{m}$ $=(s+a-b)x_{1}^{a}\partial_{2}^{b}s^{m}.$
For $\ell=b=0$, the actions of $L_{h}$ and $R_{h}$ are as in (3.82)-(3.83), and the actions of $L_{e}$ and $L_{f}$ are given by
(3.84) $L_{e}\cdot x_{1}^{k}s^{m}$ $=-kx_{1}^{k-1}x_{2}s^{m};$
(3.85) $L_{f}\cdot x_{1}^{k}s^{m}$ $=-x_{1}^{k+1}\partial_{2}s^{m}.$
As in 3.4.2, this $\widetilde{\mathcal{U}}$-module has an $n$-step filtration by powers of $s$, and decomposes into a direct sum of $\widetilde{\mathcal{U}}$-submodules, each spanned by the set of monomials $\{x_{1}^{k}x_{2}^{\ell}s^{m}\}$ and $\{x_{1}^{a}\partial_{2}^{b}s^{m}\}$ such that $k+\ell=a-b$ is a fixed integer. For $k+\ell=a-b=\lambda$, this submodule is isomorphic to a deformed Verma module of highest weight $\lambda$. We illustrate the module corresponding to $\lambda=0$ in Figure 4
Figure 4. Deformed Verma modules arise as global sections of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$.
3.4.4. Step 4: Image of the canonical map
Set $I_{U}=\mathcal{A}_{U}\langle\partial_{1},x_{2}\partial_{2}-s\rangle$ and $I_{\widetilde{X}}=\mathcal{A}_{\widetilde{X}}\langle\partial_{1},x_{2}\partial_{2}-s\rangle$ to be the left ideals generated by the operators $\partial_{1}$ and $x_{2}\partial_{2}-s$ in $\mathcal{A}_{U}$ and $\mathcal{A}_{\widetilde{X}}$, respectively. The canonical map between !!!- and $+$-pushforward is given by
$j_{!}f^{s}\mathcal{O}_{U}^{(n)}=\mathcal{A}_{\widetilde{X}}/I_{\widetilde{X}}$ $\xrightarrow{\mathrm{can}}\mathcal{A}_{U}/I_{U}=j_{+}f^{s}\mathcal{O}_{U}^{(n)}$
$1I_{\widetilde{X}}$ $\longmapsto 1I_{U}$
Since $1\mathcal{A}_{\widetilde{X}}$ generates $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ as a $\mathcal{A}_{\widetilde{X}}$-module, its image completely determines the morphism $\mathrm{can}$. On the monomial basis elements $x_{1}^{k}x_{2}^{\ell}s^{m}$ and $x_{1}^{k}\partial_{2}^{\ell}s^{m}$ of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$, the canonical map acts by
(3.86) $x_{1}^{k}x_{2}^{\ell}s^{m}\xmapsto{\mathrm{can}}x_{1}^{k}x_{2}^{\ell}s^{m}\text{ and }x_{1}^{a}\partial_{2}^{b}s^{m}\xmapsto{\mathrm{can}}s(s-1)\cdots(s-b+1)x_{1}^{a}x_{2}^{-b}s^{m}$
for $b>1$. For $b=1$, $x_{1}^{a}\partial_{2}s^{m}\xmapsto{\mathrm{can}}sx_{1}^{a}x_{2}^{-1}$.
The image of the morphism $\mathrm{can}$ is the $\mathcal{A}_{\widetilde{X}}$-submodule
(3.87) $\operatorname{im}(\mathrm{can})=\mathcal{A}_{\widetilde{X}}/I_{U}\subset\mathcal{A}_{U}/I_{U}.$
In the description of the global sections of $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ in (3.62), the global sections of $\operatorname{im}(\mathrm{can})$ can be identified with
(3.88) $\left(\mathbb{C}[x_{1},x_{2}]\otimes\mathbb{C}[s]/s^{n}+\mathbb{C}[x_{1},x_{2},x_{2}^{-1}]\otimes s\mathbb{C}[s]/s^{n}\right)t^{s}.$
3.4.5. Step 5: The maximal extension
Composing the canonical map $\mathrm{can}$ with $s$ gives
[…]
The global sections of the image of $s^{1}(n)$ (as a submodule of (3.62)) are
(3.90) $\Gamma(\widetilde{X},\operatorname{im}s^{1}(n))\simeq\left(\mathbb{C}[x_{1},x_{2}]\otimes s\mathbb{C}[s]/s^{n}+\mathbb{C}[x_{1},x_{2},x_{2}^{-1}]\otimes s^{2}\mathbb{C}[s]/s^{n}\right)t^{s}.$
This gives us an explicit description of $\Xi_{\rho}\mathcal{O}_{U}=\operatorname{coker}s^{1}(n)$:
(3.91) $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$ $=\left(\mathbb{C}[x_{1},x_{2},x_{2}^{-1}]\otimes\mathbb{C}[s]/s^{n}\right)t^{s}/\Gamma(\widetilde{X},\operatorname{im}s^{1}(n))$
(3.92) $=\left(\mathbb{C}[x_{1},x_{2},x_{2}^{-1}]\otimes\mathbb{C}[s]/s^{2}\right)t^{s}/\left(\mathbb{C}[x_{1},x_{2}]\otimes s\mathbb{C}[s]/s^{2}\right)t^{s}$
A caricature of the $\Gamma(\widetilde{X},\mathcal{A}_{\widetilde{X}})$-module (3.92) is illustrated as in Figure 5. It has two layers, corresponding to the two non-zero powers of $s$, and action by $s$ moves layers up. As vector spaces, the bottom layer is isomorphic to $\mathbb{C}[x_{1},x_{2},x_{2}^{-1}]$ and the top layer to $sx_{2}^{-1}\mathbb{C}[x_{1},x_{2}^{-1}]$.
Figure 5. Caricature of the maximal extension $\Xi_{\rho}\mathcal{O}_{U}$.
Our final step is to examine the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$. The module (3.92) has a basis given by monomials $x_{1}^{k}x_{2}^{\ell}t^{s}$ for $k\in\mathbb{Z}_{\geq 0}$ and $\ell\in\mathbb{Z}$ and $x_{1}^{k}x_{2}^{\ell}st^{s}$ for $k\in\mathbb{Z}_{\geq 0}$ and $\ell\in\mathbb{Z}_{<0}$. The actions of the operators $L_{e},L_{f},L_{h}$ and $R_{h}$ (3.12) on these monomials are given by applying the formulas (3.66)- (3.69) and taking the image of the resulting monomials in the quotient (3.92).
The $\widetilde{\mathcal{U}}$-module $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$ splits into a direct sum of submodules spanned by monomials $x_{1}^{k}x_{2}^{\ell}t^{s}$ and $x_{1}^{k}x_{2}^{\ell}st^{s}$ such that $k+\ell$ is a fixed integer. We illustrate the submodule for $k+\ell=0$ in Figure 6888For clarity, we drop the generator $t^{s}$ from our notation in Figure 6.. If $\lambda\geq 0$, the submodule corresponding to the integer $\lambda=k+\ell$ has the Verma module of highest weight $\lambda$ as a submodule, and the dual Verma module corresponding to $\lambda$ as a quotient. As a $\mathcal{U}(\mathfrak{g})$-module, it is isomorphic to the big projective module [^9] The big projective module is the projective cover of the irreducible highest weight module $L(w_{0}\lambda)$, where $w_{0}$ is the longest element of the Weyl group. It is the longest indecomposable projective object in the block $\mathcal{O}_{\lambda}$ of category $\mathcal{O}$ [Hum08, §3.12]. $P(w_{0}\lambda)$ in the corresponding block of category $\mathcal{O}$.
Figure 6. Big projective modules arise as global sections of slices of $\Xi_{\rho}\mathcal{O}_{U}$.
3.5. The monodromy filtration and the geometric Jantzen filtration
The maximal extension $\Xi_{\rho}\mathcal{O}_{U}$ naturally comes equipped with a nilpotent endomorphism $s$, giving it a corresponding monodromy filtration. This is the source of the geometric Jantzen filtrations on $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$. In this section, we use the monodromy filtration on $\Xi_{\rho}\mathcal{O}_{U}$ to compute the geometric Jantzen filtration on $j_{!}\mathcal{O}_{U}$. Using the computations of Section 3.4, we then describe the corresponding $\widetilde{\mathcal{U}}$-module filtration on global sections.
We begin by recalling monodromy filtrations in abelian categories, following [Del80, §1.6]. Given an object $A$ in an abelian category $\mathcal{A}$ and a nilpotent endomorphism $s:A\rightarrow A$, it follows from the Jacobson-Morosov theorem [Del80, Proposition 1.6.1] that there exists a unique increasing exhaustive filtration $\mu^{\bullet}$ on $A$ such that $s\mu^{n}\subset\mu^{n-2}$, and for $k\in\mathbb{N}$, $s^{k}$ induces an isomorphism $\mathrm{gr}_{\mu}^{k}A\simeq\mathrm{gr}_{\mu}^{-k}A.$ This unique filtration is called the monodromy filtration of $A$. The monodromy filtration $\mu^{\bullet}$ induces filtrations $J_{!}^{\bullet}$ on $\ker s$ and $J_{+}^{\bullet}$ on $\operatorname{coker}s$. These sub and quotient filtrations can be described explicitly in terms of powers of $s$. Namely, if we set
(3.93) $\mathscr{K}^{p}A:=\begin{cases}\ker s^{p+1}&\text{for }p\geq 0;\\ 0,&\text{for }p<0\end{cases}$
to be the kernel filtration of $A$ and
(3.94) $\mathscr{I}^{q}A:=\begin{cases}\operatorname{im}s^{q}&\text{for }q>0;\\ A&\text{for }q\leq 0,\end{cases}$
to be the image filtration of $A$, then it is shown in [Del80, §1.6] that
(3.95) $J_{!}^{i}=\ker s\cap\mathscr{I}^{-i}\text{ and }J_{+}^{i}=(\mathscr{K}^{i}+\operatorname{im}s)/\operatorname{im}s.$
In the setting of holonomic $\mathcal{D}$-modules, the filtrations $J_{!}^{\bullet}$ and $J_{+}^{\bullet}$ define the geometric Jantzen filtrations.
Definition 3.9.
Let $Y$ be a smooth variety, $f$ a regular function on $Y$ and $U=f^{-1}(\mathbb{A}^{1}-\{0\})$ as in (3.50). For a holonomic $\mathcal{D}_{U}$-module $\mathcal{M}_{U}$, recall that $j_{!}\mathcal{M}_{U}=\ker(s:\Xi_{f}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U})$ and $j_{+}\mathcal{M}_{U}=\operatorname{coker}(s:\Xi_{f}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U})$ [BB93, Lemma 4.2.1]. The filtrations $J_{!}^{\bullet}$ of $j_{!}\mathcal{M}_{U}$ and $J_{+}^{\bullet}$ of $j_{+}\mathcal{M}_{U}$ are called the geometric Jantzen filtrations.
Now we return to our running example. The monodromy filtration $\mu^{\bullet}$ on $\Xi_{\rho}\mathcal{O}_{U}$ is
(3.96) $\mu^{-2}=0\subset\mu^{-1}=\operatorname{im}s\subset\mu^{0}=j_{!}\mathcal{O}_{U}\subset\mu^{1}=\Xi_{\rho}\mathcal{O}_{U}.$
Restricting this to $\ker s=j_{!}\mathcal{O}_{U}$, we obtain the geometric Jantzen filtration of $j_{!}\mathcal{O}_{U}$:
(3.97) $0\subset\operatorname{im}s\subset j_{!}\mathcal{O}_{U}.$
The induced filtration on $\operatorname{coker}s=j_{+}\mathcal{O}_{U}$ gives the geometric Jantzen filtration on $j_{+}\mathcal{O}_{U}$:
(3.98) $0\subset\ker s/\operatorname{im}s\subset j_{+}\mathcal{O}_{U}.$
Remark 3.10.
(Deformation direction) There are other choices of regular functions on $\widetilde{X}$ which we could have used in the construction of these filtrations. In particular, if $\lambda\in\mathfrak{h}^{*}$ is dominant and integral such that $\lambda(h)=n$ for $n\in\mathbb{Z}_{\geq 0}$, then the function $f_{\lambda}:(x_{1},x_{2})\mapsto x_{2}^{n}$ can be used to define an intermediate extension functor $\Xi_{f_{\lambda}}$ and corresponding Jantzen filtrations. For $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{C})$, all choices lead to the same filtration. However, for other Lie algebras, Williamson has shown that this is not the case [Wil16].
Using the computations in Section 3.4, we can examine the $\widetilde{\mathcal{U}}$-module filtrations that we obtain on global sections. Recall that $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$ decomposes into a direct sum of submodules spanned by monomials $x_{1}^{k}x_{2}^{\ell}t^{s}$ and $x_{1}^{k}x_{2}^{\ell}st^{s}$ such that $k+\ell$ is a fixed non-negative integer. Figure 6 illustrates the submodule corresponding to $k+\ell=0$. Looking at this figure, it is clear that $\ker s=\mathrm{span}\{x_{1}^{k}x_{2}^{\ell}st^{s},t^{s}\}$ is isomorphic to the Verma module of highest weight 00, and $\operatorname{coker}s=\mathrm{span}\{x_{1}^{k}x_{2}^{\ell}t^{s}\}$ is isomorphic to the corresponding dual Verma module. Moreover, the global sections of the monodromy filtration on $\Xi_{\rho}\mathcal{O}_{U}$ restricted to the submodule corresponding to $k+\ell=\lambda$ is the composition series of the corresponding big projective module $P(w_{0}\lambda)$ when $\lambda\geq 0$. This is illustrated in Figure 7 for $\lambda=0$. We conclude that the filtrations on the Verma module $M(\lambda)$ and dual Verma module $I(\lambda)$ obtained by taking global sections of the geometric Jantzen filtrations are the composition series [^10] This is an $\mathfrak{sl}_{2}(\mathbb{C})$ phenomenon. For larger groups this procedure will yield a filtration different from the composition series..
Figure 7. Global sections of the monodromy filtration on $\Xi_{\rho}\mathcal{O}_{U}$ are the composition series of the big projective module.
3.6. Relation to the algebraic Jantzen filtration
The geometric Jantzen filtrations described above have an algebraic analogue, due to Jantzen [Jan79]. In this section, we recall the construction of the algebraic Jantzen filtration of a Verma module, then explain its relation with the geometric construction in Section 3.5.
3.6.1. The algebraic Jantzen filtration
We follow [Soe08]. Another nice reference for Jantzen filtrations is [IK11].
Let $\mathfrak{g}$ be a complex semisimple Lie algebra, $\mathfrak{b}$ a fixed Borel subalgebra, $\mathfrak{n}=[\mathfrak{b},\mathfrak{b}]$ the nilpotent radical of $\mathfrak{b}$, and $\mathfrak{h}$ a Cartan subalgebra so that $\mathfrak{b}=\mathfrak{h}\oplus\mathfrak{n}$. Denote by $\bar{\mathfrak{b}}$ the opposite Borel subalgebra to $\mathfrak{b}$. For $\lambda\in\mathfrak{h}^{*}$, denote the Verma module of highest weight $\lambda$ by
(3.99) $M(\lambda):=\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{U}(\mathfrak{b})}\mathbb{C}_{\lambda}.$
Denote by
(3.100) $I(\lambda):=\operatorname{Hom}_{\mathcal{U}(\mathfrak{b})}(\mathcal{U}(\mathfrak{g}),\mathbb{C}_{\lambda})$
the corresponding dual Verma module.
Set $T=\mathcal{O}(\mathbb{C}\rho)$ to be the ring of regular functions on the line $\mathbb{C}\rho\subset\mathfrak{h}^{*}$, where $\rho$ is the half sum of positive roots in the root system determined by $\mathfrak{b}$. A choice of linear functional $s:\mathbb{C}\rho\rightarrow\mathbb{C}$ gives an isomorphism $T\simeq\mathbb{C}[s]$. Fix such an identification. Set $A:=T_{(s)}$ to be the local $\mathbb{C}$-algebra obtained from $T$ by inverting all polynomials not divisible by $s$, and
(3.101) $\varphi:\mathcal{O}(\mathfrak{h}^{*})\rightarrow A$
to be the composition of the restriction map $\mathcal{O}(\mathfrak{h}^{*})\rightarrow T$ with the inclusion $T\hookrightarrow A$. Note that under the identification $\mathcal{U}(\mathfrak{h})\simeq\mathcal{O}(\mathfrak{h}^{*})$, $\varphi(\mathfrak{h})\subseteq(s)$, the unique maximal ideal of $A$.
For $\lambda\in\mathfrak{h}^{*}$, the deformed Verma module corresponding to $\lambda$ is the $(\mathfrak{g},A)$-bimodule
(3.102) $M_{A}(\lambda):=\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{U}(\mathfrak{b})}A_{\lambda},$
where the $\mathcal{U}(\mathfrak{b})$-module structure on $A_{\lambda}$ is given by extending the $\mathfrak{h}$-action
(3.103) $h\cdot a=(\lambda+\varphi)(h)a$
trivially to $\mathfrak{b}$. Here $h\in\mathfrak{h}$, $a\in A$, and $\varphi$ is as in (3.101).
The deformed dual Verma module corresponding to $\lambda$ is
(3.104) $I_{A}(\lambda):=\operatorname{Hom}_{\mathcal{U}(\bar{\mathfrak{b}})}(\mathcal{U}(\mathfrak{g}),A_{\lambda}).$
There is a canonical isomorphism [Soe08, Prop 2.12]
(3.105) $\operatorname{Hom}_{(\mathfrak{g},A)\mathrm{-mod}}(M_{A}(\lambda),I_{A}(\lambda))\simeq A.$
Under this isomorphism, $1\in A$ distinguishes a canonical $(\mathfrak{g},A)$-bimodule homomorphism [^11] This canonical map can be explicitly constructed using the unique nondegenerate $A$-valued contravariant form on $M_{A}(\lambda)$. This is how the filtration was originally defined in the work of Jantzen [Jan79].
(3.106) $\psi_{\lambda}:M_{A}(\lambda)\rightarrow I_{A}(\lambda).$
For any $A$-module $M$, there is a descending $A$-module filtration $M^{i}:=s^{i}M$ with associated graded $gr^{i}M=M^{i}/M^{i+1}$. Hence there is a reduction map
(3.107) $\mathrm{red}:M\rightarrow gr^{0}M=M/sM.$
For $M_{A}(\lambda)$ and $I_{A}(\lambda)$, the layers of this filtration are $\mathfrak{g}$-stable, so we obtain surjective $\mathfrak{g}$-module homomorphisms
(3.108) $\mathrm{red}:M_{A}(\lambda)\rightarrow M(\lambda)=gr^{0}M_{A}(\lambda)\text{ and }\mathrm{red}:I_{A}(\lambda)\rightarrow I(\lambda)=gr^{0}I_{A}(\lambda).$
Pulling back the filtration above along the canonical map $\psi_{\lambda}$ (3.106) gives a $(\mathfrak{g},A)$-bimodule filtration of $M_{A}(\lambda)$.
Definition 3.11.
The algebraic Jantzen filtration of $M_{A}(\lambda)$ is the $(\mathfrak{g},A)$-bimodule filtration
$M_{A}(\lambda)^{i}:=\left\{m\in M_{A}(\lambda)\mid\psi_{\lambda}(m)\in s^{i}I_{A}(\lambda)\right\},$
where $\psi_{\lambda}$ is the canonical map (3.106). By applying the reduction map (3.108) to the filtration layers, we obtain a filtration $M(\lambda)^{\bullet}$ of $M(\lambda)$.
3.6.2. Relationship between algebraic and geometric Jantzen filtrations
Though the constructions seem quite different at first glance, the geometric Jantzen filtration in Section 3.5 aligns with the algebraic Jantzen filtration described in Section 3.6.1 under the global sections functor. In this final section, we illustrate this relationship through our running example.
Recall the canonical map (3.53)
$\mathrm{can}:j_{!}f^{s}\mathcal{O}_{U}^{(n)}\rightarrow j_{+}f^{s}\mathcal{O}_{U}^{(n)}.$
As illustrated in Figures 3 and 4 , the global sections of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ and $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ decompose into direct sums of deformed dual Verma and Verma modules, respectively. [^12] To be more precise, the submodules of $\Gamma(\widetilde{X},j_{!}f^{s}\mathcal{O}_{U}^{(n)})$ and $\Gamma(\widetilde{X},j_{+}f^{s}\mathcal{O}_{U}^{(n)})$ corresponding to an integer $\lambda$ are truncated versions of $M_{A}(\lambda)$ (3.104) and $I_{A}(\lambda)$ (3.104) obtained by taking a quotient so that $s^{n}=0$.. The global sections of $\mathrm{can}$ are the direct sum of $\psi_{\lambda}$ (3.106) for all integral $\lambda$. There are two natural filtrations of $j_{!}\mathcal{O}_{U}$ which we have described using this set-up.
Filtration 1: (algebraic Jantzen filtration)
We obtain a filtration of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ by pulling back the “powers of s” filtration along $\mathrm{can}$. This induces a filtration on the quotient
(3.109) $j_{!}(\mathcal{O}_{U})\simeq j_{!}f^{s}\mathcal{O}_{U}^{(n)}/sj_{!}f^{s}\mathcal{O}_{U}^{(n)}.$
This is exactly the $\mathcal{D}$-module analogue of the algebraic Jantzen filtration described in Section 3.6.1. On global sections, it is the filtration
(3.110) $F^{i}\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})=\left\{v\in\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})\mid\Gamma(\mathrm{can})(v)\in s^{i}\Gamma(\widetilde{X},j_{+}f^{s}\mathcal{O}_{U}^{(n)})\right\}.$
Filtration 2: (geometric Jantzen filtration)
There is a unique monodromy filtration on $\Xi_{\rho}\mathcal{O}_{U}=\operatorname{coker}(s\circ\mathrm{can})$. Restricting this to the kernel of $s$, we obtain a filtration on
(3.111) $j_{!}\mathcal{O}_{U}\simeq\ker(s:\Xi_{\rho}\mathcal{O}_{U}\rightarrow\Xi_{\rho}\mathcal{O}_{U}).$
This is the geometric Jantzen filtration. It can be realized explicitly in terms of the image of powers of $s$ using 3.95. On global sections, this gives
(3.112) $G^{i}\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})=\left\{w\in\ker(s\circlearrowright\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U}))\mid w\in s^{i}\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})\right\}.$
It is helpful to see these filtrations in a picture. Figure 8 illustrates the set-up when restricted to the submodule corresponding to $\lambda=0$.
Figure 8. Relationship between the algebraic and geometric Jantzen filtrations
The map $\mathrm{can}$ is described on basis elements in (3.86). Computing these actions for $\lambda=0$, we see in Figure 8 that $\mathrm{can}$ fixes the right-most column and sends any other monomial on the left to a linear combination of monomials directly above the corresponding monomial on the right. The image of $s_{1}(n)=s\circ\mathrm{can}$ (3.89) is highlighted in grey. The quotient by this image is the maximal extension, which is outlined in the red box. The quotient 3.109 is highlighted in blue in the left hand module, and the submodule 3.111 is highlighted in blue in the right hand module.
We see that there are two copies of $j_{!}\mathcal{O}_{U}$ in this set-up: one as a quotient of the left-hand module $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$, and one as a submodule of a quotient of the right-hand module $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$. These two copies can be naturally identified as follows.
Because the submodule $sj_{!}f^{s}\mathcal{O}_{U}^{(n)}$ is in the kernel of the composition of $\mathrm{can}$ with the quotient $j_{!}f^{s}\mathcal{O}_{U}^{(n)}\rightarrow\operatorname{coker}(s\circ\mathrm{can})=\Xi_{\rho}\mathcal{O}_{U}$, the map $\mathrm{can}$ descends to a map on the quotient:
(3.113) $\overline{\mathrm{can}}:j_{!}\mathcal{O}_{U}\simeq j_{!}f^{s}\mathcal{O}_{U}^{(n)}/sj_{!}f^{s}\mathcal{O}_{U}^{(n)}\rightarrow\Xi_{\rho}\mathcal{O}_{U}.$
By construction, the map $\overline{\mathrm{can}}$ is injective. Its image is exactly $\ker(s:\Xi_{\rho}\mathcal{O}_{U}\rightarrow\Xi_{\rho}\mathcal{O}_{U})$. This is immediately apparent in Figure 8. Hence $\overline{\mathrm{can}}$ provides an explicit isomorphism which can be used to identify the two copies of $j_{!}\mathcal{O}_{U}$. Under this identification, the algebraic Jantzen filtration 3.110 and the geometric Jantzen filtration 3.112 clearly align.